A rectangular plot of land is to be enclosed by fencing. One side is along a river and so needs no fence. If the total fencing available is 1400 meters, find the dimensions of the plot to have the maximum area. (Assume that the length is greater than or equal to the width.)

Length = ? meters

Width = ? meters
Answer #1

since one side is river, the rectangles fence Perimeter will be L+zw=1400 Hence L = 1400-20. Area = Length x width. substitut